# Screen Reader Users in the Vibe Coding Era: Adaptation, Empowerment, and New Accessibility Landscape

### v1

@@ 3. Methodology
During participant recruitment, we found that most users had limited familiarity and experience with IDE-integrated AI code assistants. Since a single interview would not provide sufficient insight into their practical use and the challenges they face, we therefore conducted a two-week, three-phase study (see Figure 2). The first phase (subsection 3.3) comprised participants completing a GitHub Copilot tutorial, working on a predefined programming task with Copilot’s assistance, and participating in a semi-structured interview to share their initial impressions. In the second phase (subsection 3.4), participants incorporated Copilot into their daily programming work for two weeks, documenting significant experiences (both positive and negative) using a diary template. The final phase (subsection 3.5) consisted of a follow-up study examining participants’ real-world experiences with code assistants. This mixed-methods approach captured both in-depth, task-based insights and longitudinal, in-situ usage patterns. This study was approved by our Institutional Review Board (IRB).
Figure 2. Overview of the three-phase study design. The study included an initial study with 16 participants, a two-week phase where participants used Copilot in regular programming and record their experiences (in total of 31 recorded cases collected), and a follow-up study.
@@ 3.1. Copilot-assisted Programming
GitHub Copilot (GitHub, 2025a), released in 2021, is a widely recognized AI code assistant, with numerous studies examining how humans interact with and are impacted by it (Barke et al., 2023; Flores-Saviaga et al., 2025; Mozannar et al., 2024a). In this study, we selected Copilot within Visual Studio Code (VS Code) because of its widespread adoption, advanced feature set, and strong accessibility support (GitHub, 2025b). Copilot’s core features include code completion, inline chat, and a dedicated chat panel (see Figure 3). Code completion (Figure 3 (a)) provides real-time suggestions as users type. Inline chat (Figure 3 (b)) enables direct conversational interactions from the editor or terminal, allowing users to request code explanations or modifications for selected code snippets. The chat panel (Figure 3 (c)) serves as an integrated conversational interface, where users can pose programming questions, request code suggestions, and seek explanations.
The chat panel features three distinct modes: (1) Ask mode, for general programming queries and code generation; (2) Edit mode, for automatic application of Copilot-generated changes to one or more files in the workspace; and (3) Agent mode, where Copilot autonomously plans and executes multi-step tasks, including searching for relevant files, making code edits, and running terminal commands, iteratively refining its outputs as needed. At the bottom of the chat panel, users can send their requests and attach additional context—such as files, terminal outputs, or other relevant information—and switch between different modes and models to customize their interactions. Users can review and modify Copilot’s code changes directly within the editor (Figure 3 (1)) and navigate the message list (Figure 3 (2)). For accessibility, Copilot also offers features such as automatic playback of generated AI responses and an accessible view (Figure 3 (3)), which displays responses in a plain-text format, making it easier for screen reader users to access the information.
Figure 3. Overview of GitHub Copilot features. (a) Code Completion provides real-time code suggestions as users type. (b) Inline Chat enables seamless conversational interaction with the AI directly within the editor or terminal. (c) Chat Panel supports multi-turn conversations with three modes (i.e., Ask mode, Edit mode, and Agent mode). Users can review Copilot responses in three ways: (1) directly reviewing and editing generated code in the editor, (2) navigating the message list in the chat panel, and (3) accessing responses in the accessible view.
@@ 3.2. Participants
We recruited 16 screen reader users (2 female, 14 male) who had experience using AI for programming support. All participants used screen readers (such as NVDA (NV Access, 2025)) as their primary method of accessing development environments. 13 participants were totally blind, while 3 had low vision; of these, P8 and P9 could perceive the screen focus location. Among these participants, 7 had prior experience with Copilot, though most had only used the code completion feature and not extensively; only P12 had used the Ask mode. P10 and P14 had experience with Cursor (Cursor, 2025a) (both Ask mode and Agent mode), and P16 had used Cline (Cline, 2025) for code completion, Ask mode, and Edit mode. See Table 1 for a summary of participant demographics. Each participant received \$70 USD (or local equivalent) as compensation for their time and effort, or \$55 USD if they did not participate in the follow-up, plus an additional \$10 USD to cover a one-month GitHub Copilot subscription.
Table 1. Participants’ demographic and study-related information. Status denotes the participant’s visual status. Exp. indicates years of programming experience. Freq. denotes coding frequency, categorized as: Daily (multiple times per day), Weekly (multiple times per week), or Monthly (multiple times per month). Lang. represents the programming language used during the task. Res. indicates the number of completed requirements out of the total for each task.
@@ 3.3. Initial Study
Each study session was conducted remotely, with participants using their own familiar devices and screen readers. Before the study, participants installed the GitHub Copilot extension and set up a suitable Python or Node.js environment. We provided a step-by-step preparation guide and an accessible Copilot user manual, including feature overview, usage instructions, and keyboard shortcuts. Reviewing these materials was optional. Each session lasted 110-140 minutes, depending on how quickly participants became comfortable with Copilot. With participants’ consent, we recorded their shared screen activity and audio for analysis.
The initial user study consisted of four main steps: (1) A background interview about participants’ programming experience and AI code assistant usage. (2) A tailored Copilot tutorial based on prior experience, focusing on the introduction of key features and how to use them with screen readers, was provided. Participants were encouraged to try features they had not previously used, such as inline chat, mode switching, context setting, and different output reviewing methods. After addressing the questions, we proceeded with the coding tasks. (3) A programming task randomly chosen from four options (Table 2), with approximately one hour for completion. If unfinished, the task ended with the participant’s consent. Participants were asked to enable VS Code’s screencast mode to make their keyboard input visible, and were encouraged to think aloud—describing their thoughts, strategies, and decision-making processes while interacting with Copilot. They could use Copilot as usual or ignore it if preferred. The interviewer provided assistance and clarification as needed, particularly for accessibility barriers and Copilot usage challenges. (4) Participants rated their Copilot experience on a 7-point Likert scale (Appendix Table C), followed by a semi-structured interview about their usage patterns, challenges, and observations (Appendix Table A). Participants with experience using other code assistants shared their comparative perspectives.
@@ 3.3.1. Task
We selected four common collaborative programming tasks (see Table 2) representing a broad spectrum of real-world software development activities. These tasks required participants to implement new features based on the existing codebase and requirements (T1, T2), analyze and format data (T3), and debug others’ code (T4). In summary, these tasks covered requirements analysis, code comprehension, testing, debugging, and documentation.

### v2

@@ 3. Methodology
During participant recruitment, we found that most users had limited familiarity and experience with IDE-integrated AI code assistants. Since a single interview would not provide sufficient insight into their practical use and the challenges they face, we therefore conducted a two-week, three-phase study (see Figure 1). The study was conducted remotely between April and May 2025, with participants using their own familiar devices and screen readers. It was approved by our Institutional Review Board (IRB).
Figure 1. Overview of the three-phase study design. The study included an initial study with 16 participants, a two-week phase where participants used Copilot in regular programming and recorded their experiences (31 cases collected), and a follow-up study.:
@@ 3.1. Copilot-assisted Programming
GitHub Copilot (githubcopilot) is a widely recognized AI code assistant, with numerous studies examining how humans interact with and are impacted by it (barke2023grounded; flores2025impact; mozannar2024reading). We selected Copilot within Visual Studio Code (VS Code) because of its widespread adoption, advanced feature set, and strong accessibility support (github2025copilotaccessibility). Copilot’s core features include code completion, inline chat, and a dedicated chat panel (see Figure 2). Code completion (Figure 2 (a)) provides real-time suggestions as users type. Inline chat (Figure 2 (b)) allows users to conversationally interact with selected code snippets directly from the editor or terminal, requesting explanations or modifications. The chat panel (Figure 2 (c)) serves as an integrated conversational interface, where users can pose questions, request code suggestions, and seek explanations.
The chat panel features three modes: (1) Ask mode, for general programming queries and code generation; (2) Edit mode, for automatic application of Copilot-generated changes to one or more files in the workspace; and (3) Agent mode, where Copilot autonomously plans and executes multi-step tasks, including searching for relevant files, making code edits, running terminal commands, and iteratively refining outputs when errors occur. At the bottom of the chat panel, users can send their requests and attach context—such as files or terminal outputs—and switch between different modes and models to customize their interactions. GitHub Copilot provides four primary ways for users to review generated responses: (1) displaying code changes directly in the code editor (Figure 2 (1)); (2) structured presentation of messages and code changes in a message list (Figure 2 (2)); (3) presenting responses in a plain-text accessible view (Figure 2 (3)); and (4) automatic playback of responses via screen reader upon generation.
Figure 2. Overview of GitHub Copilot features. (a) Code Completion provides real-time code suggestions as users type. (b) Inline Chat enables seamless conversational interaction directly within the editor or terminal. (c) Chat Panel supports multi-turn conversations with three modes (i.e., Ask mode, Edit mode, and Agent mode). Users can review responses in three ways: (1) reviewing and editing generated code in the editor, (2) navigating the message list in the chat panel, and (3) accessing responses in the accessible view.:
@@ 3.2. Participants
Our study involved 16 screen reader users (2 female, 14 male) who had experience using AI for programming support. Participants were recruited via professional networks, community outreach, and snowball sampling. We screened 25 individuals based on the following criteria: at least 18 years old; primarily using screen readers (e.g., NVDA (nvda)) to program in VS Code; having programming experience in Python or JavaScript; and aiming to achieve diversity across multiple demographic and experiential factors (see Table 1). 13 participants were totally blind, while 3 were low vision; of these, P8 and P9 had limited perception of the current focus location on the screen. Among these participants, 7 had prior experience with GitHub Copilot, though most had only used the code completion feature and not extensively; only P12 had used the Ask mode. P10 and P14 had experience with Cursor (cursor) (both Ask mode and Agent mode), and P16 had used Cline (cline) for code completion, Ask mode, and Edit mode. Each participant received \$70 USD (or local equivalent) as compensation for their time and effort, or \$55 USD if they did not participate in the follow-up, plus an additional \$10 USD to cover a one-month GitHub Copilot subscription.
Table 1. Participants’ demographic and study-related information. Status denotes the participant’s visual status. Exp. indicates years of programming experience. Freq. denotes the frequency of using AI code assistants, categorized as: Daily (multiple times per day), Weekly (multiple times per week), or Monthly (multiple times per month). Lang. represents the programming language used during the task. Res. indicates the number of completed requirements out of the total for each task.:
@@ 3.3. Initial Study
Before the study, participants installed the GitHub Copilot extension and set up the Python or Node.js environment. We provided optional Copilot documentation for participants to review prior to the study. Each session lasted 110-140 minutes, depending on how quickly participants became comfortable with Copilot. With participants’ consent, we recorded their shared screen activity and audio for analysis.
The initial user study consisted of four main steps: (1) A background interview about participants’ programming experience and AI code assistant usage. (2) A tailored Copilot tutorial based on participants’ prior experience, focusing on the introduction of key features and how to use them with screen readers. Participants were encouraged to try features they had not previously used. After participants’ questions were addressed, the study then continued with the coding tasks. (3) A programming task randomly chosen from four options, with approximately one hour for completion. If unfinished, the task ended with the participant’s consent. Participants were asked to enable VS Code’s screencast mode to make their keyboard input visible, and were encouraged to think aloud—describing their thoughts, strategies, and decision-making processes while interacting with Copilot. They could use Copilot as usual or ignore it if preferred. The interviewer provided assistance and clarification as needed, particularly for accessibility barriers and Copilot usage challenges. (4) Participants rated their Copilot experience on a 7-point Likert scale (Appendix Table B), followed by a semi-structured interview about their usage patterns, challenges, and observations (Appendix Table C).
@@ 3.3.1. Task
We selected four common collaborative programming tasks (Appendix Table A) representing a broad spectrum of real-world software development activities. These tasks required participants to implement new features based on the existing codebase and requirements (T1, T2), analyze and format data (T3), and debug code (T4). In summary, these tasks covered requirements analysis, code comprehension, testing, debugging, and documentation.